# Bound States at Semiconductor–Mott Insulator Interfaces

### v2

@@ IV.2 With interface perturbation
Introducing an interface perturbation results in one (or two in special cases) additional bound state(s) with $|r_{\pm}|\neq 1$. The interface perturbation must exceed a minimum value for these additional bound states to exist. It is independent of $s_{1}$ and given by
@@
with $E_{SL}=-2T/Z\pm T_{\mathbf{k}}^{\|}+V$. These additional bound states are illustrated in Fig. 10.
We note that two (rather than one) extra states emerge directly below and above one of the Hubbard bands if $V$ and $\Delta V$ align favorably. The lower state originates from the lowest solution initially inside the band being pushed outward.
@@
Figure 10: Additional bound state solutions for different interface perturbations $\Delta V=-0.25U$ and $-0.35U$. The latter one has two bound states, as the lowest standing wave solutions is pushed out of the band. $T=0.4U$ and $V=1.5U$ in the zero parallel momentum sector and $a=6$.
@@ IV.3 Non-Symmetric Case
In the case of asymmetry, where $V_{1}\neq V_{2}$ and $\Delta V_{1}\neq\Delta V_{2}$, three out of the four equations Eq. 25, 26 govern the amplitudes $A$, $B$, $C$, and $D$, with one serving as the normalization constant. The remaining equation becomes the determining equation for the energy of the bound state solutions. These solutions are also asymmetric. In the absence of a interface perturbation, the half-space with the lower on-site potential, $V_{i}$, accumulates slightly more quasi-particle probability, typically by a single-digit percentage.
@@ V Single Interface at Mott-Néel Background
At the antiferromagnetic Mott-Néel background, the interface bound states generally comprise both the even and odd solution. Choosing $\mu\leq 0$ as the semiconducting and $\mu>0$ as the Mott insulating sites, the ansatz is:
$\psi_{\mu\leq 0}=\begin{bmatrix}p_{\mu}^{0A}\\ p_{\mu}^{1A}\\ p_{\mu}^{0B}\\ p_{\mu}^{1B}\\ \end{bmatrix}=\frac{1}{2}A\lambda_{+}^{n}\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}+\frac{1}{2}B\rho_{-}^{n}\begin{bmatrix}1\\ 1\\ -1\\ -1\end{bmatrix}$ (29)
and
@@
These equations relate the interface perturbation $\Delta V$ to the amplitudes. As in the unpolarized case, there is a minimally needed strength of the additional interface perturbation for bound state to exist, but, in contrast, this threshold is constant at $\Delta V_{\mathrm{min}}\approx\frac{2T}{Z}$. Another distinct feature is that, for bound states with an energy below the semiconducting band, only a positive interface perturbation $\Delta V>0$ supports bound states. Even though the charge background with one electron per site is the same, the different spin backgrounds lead to distinctive behavior of the bound states.
@@ VI Conclusions
Utilizing the hierarchy of correlations alongside the Fermi-Hubbard model we derived equations for the existence of bound states at single and multiple interface systems with different spin backgrounds. We account for an additional interface perturbation. These quasi-particle bound states manifest in the semiconductor as real electrons, while in the Mott insulator they are doublons and holons. Even though we used a model Hamiltonian, realistic values for the on-site potential, the hopping strength and the Coulomb repulsion can be extracted from literature.
@@ Appendix A Hierarchy of Correlations
with the Liouville superoperators ^$\widehat{\mathcal{L}}_{\mu\nu}=\left[\hat{H}_{\mu\nu},\hat{\rho}\right]$ and ^$\widehat{\mathcal{L}}_{\mu}=\commutator{\hat{H}_{\mu}}{\hat{\rho}}$ ]. The next step is the decomposition of the density operator. Any operator whose expectation value one might be interested in is computed from a subset of lattice sites. This allows for the decomposition into on-site and correlated parts:
@@
with the symmetrized from ^$\widehat{\mathcal{L}}^{S}_{\mu\nu}=\widehat{\mathcal{L}}_{\mu\nu}+\widehat{\mathcal{L}}_{\nu\mu}$ for the on-site density operator. This time evolution contains the two-point correlator $\hat{\rho}_{\mu\nu}^{\mathrm{corr}}$. The same is done for the two-site density operator, such that we can combine these two results to find:
@@ C.1 Defining Equation
whose zeros give the bound states. $a$ and $b$ are the $+$ or --, respectively. The imaginary “bells” coincide with the Mott and semiconductor band, respectively, and are not affected by the interface perturbation. The term $-\Delta VZ/T$ shifts the real part of the defining equation up and down. In Fig. 11 the defining equation for $r_{a}=r_{+}$ and $s_{b}=s_{-}$ is shown without and with a interface perturbation together with the absolute value $|r_{a}|$ and $|s_{b}|$. Without any interface perturbation the zeros of the real part are inside the non-zero imaginary part. So, by adding the interface perturbation it follows directly that any solution is outside of the bands and has imaginary $\kappa_{\mathrm{Mott}}$ and $\kappa_{\mathrm{semi}}$. The conditions $|r_{a}|>1$ and $|s_{b}|<1$ then dictate on which side of the “bells” the solution needs to be, because the absolute value (at least for the semiconductor) is a monotone function of energy $E$ interrupted by the plateau of the band. For the Mott, this is also true, but at $E=U/2$ there the absolute value goes down to zero again.

### v3

@@ IV.2 With Interface Perturbation
Introducing an interface perturbation results in one (or in special cases two) additional bound state(s) with $e^{i\kappa_{\mathrm{Mott}}\mu}\neq 1$. In the Mott region these are not standing waves, but the wave function amplitude decays away from the interface, see Fig. 10. The interface perturbation must exceed a minimum value for these additional bound states to exist, which is given by:
@@
with $E_{SL}=\pm 2T/Z-T_{\mathbf{k}}^{\|}+V$. The $\pm$ moves the bound state below (-) or above (++) the semiconducting band. These additional bound states are illustrated in Fig. 10.
We note that two (rather than one) extra states emerge directly below and above one of the Hubbard bands if the band offset $V$ and the perturbation $\Delta V$ align favorably. The lower state originates from the lowest solution initially inside the band being pushed outward.
@@
Figure 10: Additional bound state solutions for different interface perturbations $\Delta V=-0.25U$ and 0.35U-0.35U. The latter one has two bound states, as the lowest standing wave solutions is pushed out of the band. $T=0.4U$ and $V=1.5U$ in the zero parallel momentum sector and $a=6$.
@@ IV.3 Non-Symmetric Case
The band offset of the semiconductors coupled to the Mott insulator do not necessarily need to be the same. In the case of asymmetry, where $V_{1}\neq V_{2}$ and $\Delta V_{1}\neq\Delta V_{2}$, three out of the four Eq. 27, Eq. 28 govern the amplitudes $A$, $B$, $C$, and $D$, with one serving as the normalization constant. The remaining equation becomes the determining equation for the energy of the bound state solutions. These solutions are also asymmetric. In the absence of an interface perturbation, the half-space with the lower band offset accumulates slightly more quasi-particle probability.
@@ V Single Interface at Mott-Néel Background
As discussed in Sec. II.4 and Sec. II.5, one may consider two mean-field background solutions in the Mott insulator. Previously we discussed the unpolarized Mott state without any spin ordering whereby the doublons and holons may simultaneously live on the same lattice site. This is in contrast to the bi-partite Mott-Néel state where the doublon and holon wave functions are exclusively non-zero on one of the sublattices, but remain coupled across lattice sites. In the setting chosen here, the holon wave function only lives on sublattice $A$ ($p_{\mu}^{0A}\neq 0$ in Mott), while the doublon one lives on sublattice $B$ ($p_{\mu}^{1B}\neq 0$ in Mott). In this antiferromagnetic Mott-Néel background, the interface bound states generally comprise both the even and odd solution. Choosing $\mu\leq 0$ as the semiconducting and $\mu>0$ as the Mott insulating sites, the ansatz for the quasi-particle wave function is
$\psi_{\mu\leq 0}=\begin{bmatrix}p_{\mu}^{0A}\\ p_{\mu}^{1A}\\ p_{\mu}^{0B}\\ p_{\mu}^{1B}\\ \end{bmatrix}=\frac{1}{2}A\lambda_{+}^{n}\begin{bmatrix}1\\ 0\\ 1\\ 0\end{bmatrix}+\frac{1}{2}B\rho_{-}^{n}\begin{bmatrix}1\\ 0\\ -1\\ 0\end{bmatrix}$ (32)
in the semiconducing sites. It is chosen to represent the conduction band. In the Mott insulating half-space it reads:
@@
These equations relate the interface perturbation $\Delta V$ to the amplitudes $A$ and $B$. As in the unpolarized case, there is a minimally needed strength of the additional interface perturbation for bound state to exist, but, in contrast, this threshold is constant at $\Delta V_{\mathrm{min}}\approx\frac{2T}{Z}$. Another distinct feature is that, for bound states with an energy below the semiconducting band, only a positive interface perturbation $\Delta V>0$ supports bound states. Even though the charge background with one electron per site is the same, the different spin backgrounds lead to distinctive behavior of the bound states.
@@ VI Conclusions
Utilizing the hierarchy of correlations alongside the Fermi-Hubbard model we derived equations for the existence of bound states of the doublon and holon wave function at single and multiple interface systems with different spin backgrounds. We account for an additional interface perturbation. These quasi-particle bound states manifest in the semiconductor as real electrons, while in the Mott insulator they are doublons and holons. Even though we used a model Hamiltonian, realistic values for the on-site potential, the hopping strength and the Coulomb repulsion can be extracted from literature.
@@ Appendix A Hierarchy of Correlations
with the Liouville superoperators $\widehat{\mathcal{L}}_{\mu\nu}=\left[\hat{H}_{\mu\nu},\hat{\rho}\right$] and $\widehat{\mathcal{L}}_{\mu}=\commutator{\hat{H}_{\mu}}{\hat{\rho}}$. The next step is the decomposition of the density operator. Any operator whose expectation value one might be interested in is computed from a subset of lattice sites. This allows for the decomposition into on-site and correlated parts:
@@
with the symmetrized from $\widehat{\mathcal{L}}^{S}_{\mu\nu}=\widehat{\mathcal{L}}_{\mu\nu}+\widehat{\mathcal{L}}_{\nu\mu}$ for the on-site density operator. This time evolution contains the two-point correlator $\hat{\rho}_{\mu\nu}^{\mathrm{corr}}$. The same is done for the two-site density operator, such that we can combine these two results to find:
@@ C.1 Defining Equation
whose zeros give the bound states. $a$ and $b$ are the ++ or -, respectively. The imaginary “bells” coincide with the Mott and semiconductor band, respectively, and are not affected by the interface perturbation. The term $\Delta VZ/T$ shifts the real part of the defining equation up and down. In Fig. 11 the defining equation for $r_{a}=r_{+}$ and $s_{b}=s_{-}$ is shown without and with a interface perturbation together with the absolute value $|r_{a}|$ and $|s_{b}|$. Without any interface perturbation the zeros of the real part are inside the non-zero imaginary part. So, by adding the interface perturbation it follows directly that any solution is outside of the bands and has imaginary $\kappa_{\mathrm{Mott}}$ and $\kappa_{\mathrm{semi}}$. The conditions $|r_{a}|>1$ and $|s_{b}|<1$ then dictate on which side of the “bells” the solution needs to be, because the absolute value (at least for the semiconductor) is a monotone function of energy $E$ interrupted by the plateau of the band. For the Mott, this is also true, but at $E=U/2$ there the absolute value goes down to zero again.